OpenStudy (anonymous):

8) g(n) = 3n + 2 Perform the indicated operation f (n) = 2n 2 + 5 Find g( f (2)) Seems simple, but I'm a little stuck. Maybe if you can explain it too, that'd be great.

OpenStudy (anonymous):

f (n) = 2n +5 is the second bit?

OpenStudy (anonymous):

Woops the first part and second part are supposed to be together. 8) g(n) = 3n + 2 f (n) = 2n 2 + 5 Find g( f (2))

OpenStudy (anonymous):

so one equation is g(n) + 3n +2 and the other is f(n) = 2n? i'm not getting what the 2 and the +5 are after

OpenStudy (anonymous):

g(n) = 3n + 2 f (n) = 2n^2 + 5 Find g( f (2))

OpenStudy (anonymous):

riightt do f(2) first.

OpenStudy (anonymous):

plug in 2 for in in the function f(n) = ...

OpenStudy (anonymous):

2(2)^2 + 5 2*4+5 8+5 13

OpenStudy (anonymous):

right, no you have g(13) left to finish it off

OpenStudy (anonymous):

now'

OpenStudy (anonymous):

3(13) + 2 right?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

41. Wow that was easy. Thanks!

OpenStudy (anonymous):

no problem!